OpenStudy (anonymous):

PLEASE HELP There are (7^13)^3 ⋅ 7^0 strawberries in a field. What is the total number of strawberries in the field? if you understand how to do this, please dont just give the answer, please explain how to do it

OpenStudy (anonymous):

Answer is 7^39, let me type out the work now

OpenStudy (anonymous):

7^0=1

OpenStudy (anonymous):

This leaves us with ((7^13)3)*1

OpenStudy (anonymous):

((7^13)^3)=7^39

OpenStudy (anonymous):

7^39*1=7^39

OpenStudy (anonymous):

im confused because like do i multiply 3 x 91 or do i leave 7^13 the same

OpenStudy (anonymous):

@Ilovecake you can leave if you want, this guy/girl is teaching me (:

OpenStudy (anonymous):

leave &^13 the same.

OpenStudy (anonymous):

*7^13

OpenStudy (anonymous):

ohhhh! okay! thank you soooooooooooooo much!

OpenStudy (ilovecake):

Ok I was trying to figure this out but ok

OpenStudy (anonymous):

Here, let me check this, make sure I am right

OpenStudy (anonymous):

im just gonna trust you when you say that because yesterday someone put a link under my question and it was a super scary face with scary sounds so lol

OpenStudy (anonymous):

Haha, ok. I do promise that those are legit though :P

OpenStudy (anonymous):

i trust you lol

OpenStudy (ilovecake):

Oh that link. I almost peed my pants

OpenStudy (anonymous):

ikr

OpenStudy (ilovecake):

its the anne one right?

OpenStudy (anonymous):

i was home alone and i was already scared cause i was at this haunted island yesterday and like wow. and idk it had a scary face that was like blinking

OpenStudy (ilovecake):

Yeah ik

OpenStudy (ilovecake):

http://akk.li/pics/anne.jpg it is this link right?

OpenStudy (anonymous):

im not opening that, are you crazy

OpenStudy (ilovecake):

No I was just asking if that was the same link

OpenStudy (ilovecake):

I am not opening it either

OpenStudy (anonymous):

i think so. but im not opening it

OpenStudy (ilovecake):

Want me to see?

OpenStudy (ilovecake):

ewwwwww it was!!! ugh

OpenStudy (anonymous):

well dont scare your self lol.

OpenStudy (ilovecake):

well I didn't know. I was just looking

OpenStudy (anonymous):

im gonna start sending that to people

OpenStudy (ilovecake):

LOL

OpenStudy (ilovecake):

I am post it as a question

OpenStudy (ilovecake):

cya